OpenStudy (anonymous):

William is 16 years older than Ashley. Sixteen years ago, William was 5 times as old as Ashley. How old is Ashley now?

zepdrix (zepdrix):

Let W be William's age. Let A be Ashley's age. W = A + 16 <--- William = Ashley + 16 years W - 16 = 5A <--- Minus 16 years from William (years ago) = 5 times Ashley age. I know these are a little tricky to set up :) Try to read little keywords like (is means = usually).

zepdrix (zepdrix):

So you have a system of equations, solve for the unknowns. Yay.